### v1

# Incompatibility of rank-one POVMs and uncertainty relations

Abstract.
The incompatibility between quantum measurements is a fundamental feature of quantum mechanics with profound implications for uncertainty relations and quantum information processing. In this paper, we realized that the $s$-order incompatibility of measurements by Xu (inspired by De Bièvre’s “complete incompatibility” ), can be equivalently characterized by spark of the associated measurement matrix. This provides a new linear algebraic perspective on measurement incompatibility. Leveraging this insight, we extend $s$-order incompatibility to rank-one POVMs and establish novel uncertainty relations. Furthermore, we explore the incompatibility of multiple bases and its connection to support uncertainty relations. These results may have applications in quantum cryptography, quantum state reconstruction, and quantum compressed sensing.
1. Introduction
The incompatibility of measurements is a fundamental feature of quantum mechanics, distinguishing quantum phenomena from classical physics. Recent efforts to rigorously quantify incompatibility have revealed profound connections to uncertainty relations, quantum state reconstruction, and the underlying structure of quantum computational advantage.
The notion of ‘incompatibility’ has different formulations in literature, see e.g. the survey paper [8]. Traditionally, incompatibility has been analyzed through the commutator of two quantum operators. However, recent studies have highlighted the significance of the anticommutator or Jordan product as an equally important descriptor of measurement incompatibility. This shift in perspective is motivated by the mathematical fact: for non-negative rank-one operators $A$ and $B$, the non-negativity condition $\{A,B\}\geq 0$ holds if and only if they commute, i.e., $[A,B]=0$. This equivalence reveals a deep duality between algebraic noncommutativity and geometric negativity in operator space – a connection with far-reaching implications for characterizing quantum coherence, see e.g. [5] and [10].
In a different approach, De Bièvre in [1] introduced the concept of complete incompatibility for two orthonormal bases of a Hilbert space $\mathscr{H}$, linking it to the support uncertainty inequality and the identification of non-classical quantum states whose Kirkwood–Dirac quasiprobability distributions fail to be proper probability distributions. Such states have been shown to provides advantages in quantum metrology and have important implications in weak measurements and contextuality. Xu subsequently generalized in [4] this concept to $s$-order incompatibility for a positive integer $s$, where the case $s=d+1$ recovers De Bièvre’s original definition. A key distinction of $s$-order incompatibility is that it takes only a finite set of integer values, contrasting with conventional continuous measures of incompatibility, such as those concepts based on Jordan product in [5].
We observe that the complete incompatibility or the more general $s$-order incompatibility for two orthonormal bases can be characterized using a concept from linear algebra — the spark of the matrix formed by concatenating the basis vectors. In the context of compressed sensing, the spark represents the smallest number of linearly dependent columns in a matrix, providing a concrete and computable measure of incompatibility. This insight offers a novel linear algebraic perspective on measurement incompatibility, linking it to sparsity and compressed sensing techniques.
Extending $s$-order incompatibility to general positive operator-valued measures (POVMs) via tight frame theory [12], we address the challenge of characterizing non-projective measurement incompatibility. This extension reveals that the incompatibility of POVMs is governed by the linear dependence structure of their measurement vectors, making it particularly useful where non-projective measurements are essential, such as quantum state estimation with noise resilience. Consequently, we derive a new class of uncertainty relations that emphasize the interplay between measurement incompatibility and the minimal support of quantum states across different measurement settings. Notably, we offer a novel interpretation of the Ghobber-Jaming uncertainty relation [3] through the lens of measurement incompatibility. This interpretation is further generalized to rank-one POVMs, addressing a question raised in [15] for general frames.
Furthermore, we extend our framework to systems involving multiple orthonormal bases, introducing a generalized $s$-order incompatibility criterion. Recall that variance-based uncertainty relations have been established for multiple observables associated with different bases, see e.g. [13]. Our approach enables the formulation of sharp multi-measurement support uncertainty relations. Such relations are crucial in applications requiring compatibility constraints across several observables, such as quantum error correction and multi-parameter metrology. Unlike the two-bases case, where the spark serves as a useful tool, the multi-bases setting requires a different generalization of the concepts by De Bièvre and Xu. Notably, our support uncertainty relation has sharp lower bound, in contrast to the continuous setting, where establishing sharp bounds for multiple measurements remains an open challenge.
These findings not only deepen our understanding of measurement incompatibility but also have potential applications in quantum cryptography, quantum state reconstruction, and quantum error correction. In particular, the spark-based characterization may provide new insights into designing and analyzing measurement schemes in quantum compressed sensing and quantum tomography.
Paper Organization and Notations: The remainder of this paper is structured as follows. In Section 2 establishes the equivalence between $s$-order incompatibility and the spark of measurement matrix. Section 3 extends this concept to POVMs via frame theory. Section 4 derives novel uncertainty relations within the incompatibility framework. Section 5 explores the implications of incompatibility in multi-basis settings and its connection to support uncertainty. Finally, we conclude by discussing potential applications and open questions.
Throughout the paper, the set of consecutive integers $\{1,2,\ldots,d\}$ is denoted by $[[1,d]]$. For two orthogonal bases $A$ and $B$, we denote $(A,B)$ by the matrix whose columns are given by the basis of $A$ and $B$. For a set $S$, the notion $|S|$ denotes the number of elements in $S$, ${\rm span}\{S\}$ denotes the space spanned by $S$ over the complex field $\mathbb{C}$. Let $I=\{1,2,\ldots,m\}$ and $J=\{1,2,\ldots,n\}$. For a set $E\subset I$, we will write $E^{c}$ for its complement.
2. $s$-order incompatibility and the spark of Measurement Matrices
Let $A=\{\ket{a_{j}}\}_{j=1}^{d}$ and $B=\{\ket{b_{k}}\}_{k=1}^{d}$ be two orthonormal bases for a $d$-dimensional complex Hilbert space $\mathscr{H}$. As usual to avoid the freedom |$\ket{a_{j}}\rightarrow e^{i\theta_{j}}\ket{a_{j}}$ with $\theta_{j}\in\mathbb{R}$ real number and $i=\sqrt{-1}$, we consider the sets of rank-one projectors $\overline{A}=\{\ket{a_{j}}\bra{a_{j}}\}_{j=1}^{d}$ and $\overline{B}=\{\ket{b_{k}}\bra{b_{k}}\}_{k=1}^{d}$. When $\overline{A}$ and $\overline{B}$ commute, we say $A$ and $B$ are compatible. Otherwise we say $A$ and $B$ is incompatible. De Bièvre introduced the concept of complete incompatibility in [1], which is purely algebraic but has a clear physical interpretation in terms of selective projective measurements. It has been shown that this concept closely related to minimal support uncertainty, and it also is useful to characterize the Kirkwood-Dirac nonclassicality. The incompatibility of two bases is further classified by Xu in [4] which includes De Bièvre’s complete incompatibility as a special case.
Definition 2.1 ($s$-order incompatibility [4]).
Two orthonormal bases $A$ and $B$ are $s$-order incompatible, where $s\in[[2,d+1]]$, if
  1. (1)
    For all nonempty subsets $S_{A}\subset A$ and $S_{B}\subset B$ with $|S_{A}|+|S_{B}|<s$, then
    ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}=\{0\}.$
  2. (2)
    There exist subsets $S_{A}$ and $S_{B}$ with $|S_{A}|+|S_{B}|=s$, such that
    ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}\neq\{0\}.$
On the other hand, we recall the spark of a matrix from compressed sensing in [11] by D.L. Donoho and M. Elad.
Definition 2.2 (Spark or Kruskal rank).
The spark of a $m\times n$ matrix $M$ is the smallest integer $k$ such that there exists a set of $k$ columns in $M$ which are linearly dependent.
Now, we have the following observation. The proof is omitted.
Theorem 2.3.
Two orthonormal bases $A$ and $B$ are $s$-order incompatible if and only if ${\rm spark}(A,B)=s,$ where $(A,B)$ is the matrix whose column vectors are the basis vectors.
3. Incompatibility of rank-one POVMs via tight frames
A Positive Operator-Valued Measure (POVM) is a set of non-negative Hermitian operators $\{Q_{j}\}$, not necessarily projectors, that satisfy $\sum_{j}Q_{j}={\rm\mathbb{I}}$ where ${\rm\mathbb{I}}$ is the identity operator on $\mathscr{H}$. A POVM generalizes the standard measurement, as its measurement vectors $\mu_{j}$ need not to be normalized or orthogonal. There exists a fundamental relationship between rank-one quantum measurements and the so-called tight frames [12]. Frames generalize bases and were introduced in the context of nonharmonic Fourier series by Duffin and Schaeffer.
Definition 3.1.
Let $\{\varphi_{j},1\leq j\leq n\}$ denote a set of (overcomplete) vectors in a $d$-dimensional space $\mathscr{H}$. The vectors $\{\varphi_{j}\}$ is said to form a tight frame if there exists a constant $\alpha>0$ such that for all $x\in\mathscr{H}$,
$\sum_{j=1}^{n}|\langle x,\varphi_{j}\rangle|^{2}=\alpha^{2}\|x\|^{2}.$
If $\alpha=1$, the tight frame is said to be normalized, otherwise, it is $\alpha$-scaled.
Theorem 3.2.
[12] A set of vectors $\varphi_{j}\in\mathscr{H}$ forms a $\alpha$-scaled tight frame for $\mathscr{H}$ if and only if the scaled vectors $\alpha^{-1}\varphi_{j}$ are the measurement vectors of a rank-one POVM on $\mathscr{H}$. In particular, the vectors $\varphi_{j}$ form a normalized tight frame for $\mathscr{H}$ if and only if they are the measurement vectors of a rank-one POVM on $\mathscr{H}$.
When no confusion can arise, we will use the same notation for the POVMs and the corresponding set of measurement vectors.
Definition 3.3 ($s$-order incompatibility for POVMs).
Two rank-one POVMs $A$ and $B$ are called $s$-order incompatible if ${\rm spark}(A,B)=s$, where $(A,B)$ are the measurement vector matirx.
For a pure state |$\ket{\varphi}$, let $\{\langle a_{j}|\varphi\rangle\}$ and $\{\langle b_{k}|\varphi\rangle\}$ denote the coefficients under the POVMs. Denote $n_{A}(\ket{\varphi})$ for the number of non-zero elements in $\{\langle a_{j}|\varphi\rangle\}$ and $n_{B}(\ket{\varphi})$ as the number of non-zero elements in $\{\langle b_{k}|\varphi\rangle\}$. Then the following sharp support uncertainty relation holds.
Theorem 3.4 (support uncertainty relation).
For $s$-order incompatible rank-one POVMs $A$ and $B$, Then for any non-zero pure state |$\ket{\varphi}$, we have
$n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})\geq s.$
Proof.
It follows from the property of spark, as detailed in [11]. ∎
This yields an equivalent definition of the $s$-order incompatibility for a pair of POVMs or for general tight frames.
Definition 3.5.
Suppose $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are two tight frames in $\mathscr{H}$, with frame constant $\alpha$ and $\beta$. If the integer $s$ satisfies
  1. (1)
    For any non-empty subsets $S\subseteq I$ and $J\subseteq J$, satisfying $|S|+|T|<s$, such that for any non-zero $x\in\mathscr{H}$,
    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
    do not hold simultaneously,
  2. (2)
    There exist non-empty subsets $S\subseteq I$ and $T\subseteq J$, satisfying $|S|+|T|=s$, such that there exists a non-zero $x\in\mathscr{H}$, the following
    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
    hold simultaneously.
Then $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are called $s$-order incompatible.
4. Uncertainty relations from $s$-order incompatibility
In this section, we establish new uncertainty relation based on the concept of s-order incompatibility. By leveraging the spark of measurement matrices, we derive lower bounds on the support of quantum states under different measurement settings, extending traditional uncertainty principles to a broader class of quantum measurements. Let $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ be two frames in a Hilbert space $\mathscr{H}$, where each set corresponds to a rank-one positive operator-valued measure (POVM).
Definition 4.1 (Minimal Reconstruction Number).
For frames $A$ and $B$, define
$t_{\min}=\min\{t\,|\,\forall S\subsetneq I,T\subsetneq J,|S|+|T|>t,\operatorname{span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)=\mathscr{H}\}.$
For any pair of $s$-order incompatible tight frames, the relationship between $s$ and $t_{\min}$ is given by the following theorem:
Theorem 4.2.
If two tight frames $A$ and $B$ are $s$-order incompatible, then
$s+t_{\min}=m+n.$
Proof.
We show that $t_{\min}=m+n-s$.
1) For any subsets $S\subsetneq I$ and $T\subsetneq J$ satisfying $|S|+|T|>m+n-s$, it follows that their complements $S^{c}$ and $T^{c}$ satisfy
$|S^{c}|+|T^{c}|<s.$
Since the frames $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ are $s$-order incompatible, thus by the equivalent Definition 3.5, for any non-zero vector $x\in\mathscr{H}$, the following two identities
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
cannot simultaneously hold. This implies that
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}<(\alpha+\beta)\|x\|^{2},\qquad\forall\,x\neq 0.$
It is equivalent to say, for any non-zero $x\in\mathscr{H}$,
$\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}>0,$
thus $\left(\operatorname{span}\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)^{\perp}=\{0\}$. It follows that
$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
Therefore, $t_{\min}\leq m+n-s$.
2) By the definition of $s$-order incompatibility, there exist two non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|=s$, such that for some nonzero $x$, the following hold
$\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
This means that for any $S^{c}\subsetneq I$ and $T^{c}\subsetneq J$ satisfying $|S^{c}|+|T^{c}|=m+n-s$, there exists non-zero $x\in\mathscr{H}$ such that
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=0.$
Thus $\left({\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\right)^{\perp}\neq\{0\}$, i.e.
${\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\neq\mathscr{H}.$
Hence $t_{\min}\geq m+n-s$.
Collecting all, we obtain $t_{\min}=m+n-s$, completing the proof. ∎
Remark 4.3.
The quantity $t_{\min}$ establishes a fundamental limit on the number of measurements needed to reconstruct a quantum state from two rank-one POVMs.
Theorem 4.4.
Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ in $\mathscr{H}$ be two $s$-order incompatible tight frames. Then there exists a constant $C$ such that
(4.1) $\|x\|^{2}\leq C\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right).$
for all subsets $S\subseteq I$ and $T\subseteq J$ with $|S|+|T|<s$.
Proof.
By Theorem 4.2, for any non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|<s$, we know that
$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
Thus the set $\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}$ forms a frame for $\mathscr{H}$ (by [14, Lemma 1.2]), whose lower frame bound is denoting by $C_{S,T}$. Denoting $C_{1}=\min C_{S,T}$, the desired constant can be chosen as $C=1/\min\{\alpha,\beta,C_{1}\}$. ∎
Remark 4.5.
If the columns of the matrix $M$ are normalized to unit norm, then the lower bound of its spark can be given by its dictionary coherence
${\rm spark}(M)\geq 1+\frac{1}{\mu(M)},$
where the dictionary coherence $\mu(A)$ is defined as the maximum correlation between any to columns
$\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|.$
It is seen that our uncertainty inequality provides a more essential interpretation for Ghobber-Jaming uncertainty relation in [3] in terms of the coherence $\mu(M)$.
5. Ghobber-Jaming uncertainty relation for frames
Theorem 5.1.
Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ and $\alpha_{2},\beta_{2}$, respectively. Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying $S\subseteq T$, the condition
$\frac{\mu^{2}}{\alpha_{1}}<\frac{\beta_{2}}{T}$
holds, where $\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|$. Then for all $x\in H$,
$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}.$
Proof.
Let $x=x_{1}+x_{2}$, where
$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
in which $a_{k}^{*}$ is the canonical dual frame.
$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ \leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|\right)\\ \leq\mu^{2}\alpha_{2}|T|\|x\|^{2}.\end{split}$
Thus,
$\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}\mu^{2}\mu^{2}|T|\|x\|^{2}$
This implies
$\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|&\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}\mu^{2}|T|\|x\|^{2}\right)^{\frac{1}{2}}\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|.\end{split}$
On the other hand,
$\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ \leq m^{2}M(A^{*},A)^{2}\sum_{k\in s^{c}}|\langle x,a_{k}\rangle|^{2}\end{split}$
Thus we have
$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|&\leq\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\|x\|\end{split}$
we obtain
$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2}\end{split}$
where
$C(S,T)=\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\right)\left(\max\{\frac{1}{\beta_{1}^{1/2}},\left\{\beta_{1}^{-\frac{1}{2}},\left((\frac{\beta_{2}}{\alpha_{1}\beta_{1}})^{\frac{1}{2}}+\alpha_{1}^{-\frac{1}{2}}\right)mM(A^{*},A)\right\}\right)^{-1}.$
6. Incompatibility of multiple bases
We consider the incompatibility of multiple bases in this section.
Definition 6.1.
Orthonormal bases $A_{k}=\{\ket{a^{(k)}_{j}}\}_{j=1}^{n}$ are $s$-order incompatible if:
  1. (1)
    For any nonempty $S_{k}\subset A_{k}$ with $\sum_{k=1}^{n}|S_{k}|<s$, $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}=\{0\}.$
  2. (2)
    There exist subsets $S_{k}\subset A_{k}$, such that $\sum_{k=1}^{n}|S_{k}|=s$ and $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}\neq\{0\}.$
Theorem 6.2.
For $s$-order incompatible bases $\{A_{k}\}_{k=1}^{n}$, any non-zero pure state |$\ket{\varphi}$ if and only if
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi})\geq s.$
Proof.
Let
$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi}).$
Then if $n_{\min}=s$, there exists a pure state |$\ket{\psi}$ such that
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=s,$
and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<s$.
For such |$\ket{\psi}$, there exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi})$, and |$\ket{\psi}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. The nonexistence of such |$\ket{\psi^{\prime}}$ implies that there do not exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi^{\prime}})$, and |$\ket{\psi^{\prime}}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. These two conditions precisely coincide with (1.1) and (1.2) in Definition 1. Then the claim follows. $\square$
7. Conclusion
By equating $s$-order incompatibility with the spark, we unify quantum measurement theory with compressed sensing. Key contributions include:
  • POVM Incompatibility: Generalized to non-projective measurements via tight frames.
  • Sharp Uncertainty Bounds: Derived from the spark, improving coherence-based estimates.
  • Multi-Basis Framework: Extended support uncertainty to multi-observable systems.
Future directions include applications in quantum tomography, cryptographic protocol design, and infinite-dimensional spark generalizations.